OpenStudy (anonymous):

Find the solutions of x² -2ax -b = 0 by completing the square. The answer is x = ±√b+a² + a

OpenStudy (tkhunny):

Have you considered Completing the Square? \(x^{2} - 2ax + ??? = b + ???\) Fill in the blanks.

OpenStudy (anonymous):

\[x^2-2 a x=b\] add both sides \[\left( \frac{ -2 a }{ 2 } \right)^2 ~i.e.,a^2\]